# Velocity Analysis of Moving Objects in Earth Observation Satellite Images Using Multi-Spectral Push Broom Scanning

### v1

@@ 1 Introduction
The commercialization of space presents new opportunities in the quantity and character of available data. PLC satellites have unique capabilities with higher spatial resolution ($\sim 4$ m) and a more frequent revisit rate (daily) than the NASA Landsat or ESA Sentinel satellites (10 - 15 m resolution and revisit rates of several days). The scientific use of commercial data can be challenging if proprietorship restricts information on performance and procedures and results in omissions of data such as time, one of the fundamental properties of the universe along with space, energy and mass.
@@ 2.1 Design of the study
The camera on the PLC SuperDove satellites covers eight spectral bands with a total field of view (FOV) of $\sim 35\times 21\ {\rm km}^{2}$ and each spectral band covering one eighth the total FOV (Planet Labs Corporation, 2021). With push broom scanning and a typical orbital altitude of 500 km, the satellites observe a region on the ground in all eight spectral bands within $\sim 3.2$ s. We assume that the camera operates asynchronously with the travel of the image footprint and with a constant rate over each image sufficient to obtain overlapping images in each spectral band. These smaller, single-band image strips, $\sim 35\times 2.6\ {\rm km}^{2}$, are mosaiced into larger images with approximately the same size as the total FOV. Different regions within the larger, mosaiced image then have different acquisition times, each corresponding to the acquisition time of the smaller, overlapping strip used in each particular region.
@@ 2.3 Time delays between positions
Figure 2: Track A in table 2.2. The circles show the locations of an aircraft identified from the individual spectral bands over a visual image (left) and over the sum of the absolute value of the seven differenced images (right). There is no circle for the NIR band because the signal-to-noise is too low to allow for precise localization. The location in the NIR band is between the two locations at the head of the track (top of the image). The spacing between the bottom two locations is longer than the others, not counting the gap due to the missing NIR location. This longer spacing is due to a shift in the mosaicing pattern that adds a time delay equal to the time between exposures of the camera, $\Delta t_{camera}$, to the time delay equal to the crossing time, $\Delta t_{band}$, of one of the single-band image strips that are combined into a larger single-band image.
@@ 3 Results
Table 2: Mosaic time correction parameters $\alpha_{i}$ in equation 1 for the tracks A-I
[…]
@@
[…]
1 Indicates the segment measured between the positions in spectral bands 1 and 5. The segments are arranged in time order. 2 Because the NIR position (band 7) is not used, $\alpha_{i}$ applies to the difference in time between the positions in bands 6 and 0. 3 The ellipses indicate that a segment between the positions in two bands is not defined because one or both positions can not be determined due to low S/N. In this case the parameter $\alpha$ in the previous column applies to the segment from spectral band 3 to 6.
@@
Table 3: True velocity depending on altitude for track G
[…]
@@
[…]
1 Above ground level (AGL) = aircraft geo-altitude - altitude of DFW.
@@
[…]
1 The standard deviation is a measure of the accuracy of the velocity calculated from the PLC images, specifically the variation in the velocity for each segment of the track of an aircraft.

### v2

@@ 1 Introduction
The commercialization of space presents new opportunities in the quantity and character of available data. PLC satellites have unique capabilities with higher spatial resolution ($\sim 4$ m) and a more frequent revisit rate (daily) than the NASA Landsat or ESA Sentinel satellites (10 - 15 m resolution and revisit rates of several days). The scientific use of commercial data can be challenging if proprietary restrictions on information, performance and procedures results in omissions of data such as time, one of the fundamental properties of the universe along with space, energy and mass.
@@ 2.1 Design of the study
The camera on the PLC SuperDove satellites covers eight spectral bands with a total field of view (FOV) of $\sim 35\times 21\ {\rm km}^{2}$ and each spectral band covering one eighth the total FOV (Planet Labs Corporation, 2021). With push broom scanning and a typical orbital altitude of 500 km, the satellites observe a region on the ground in all eight spectral bands within $\sim 3.2$ s. We assume that the camera operates asynchronously with the travel of the image footprint and with a constant rate over each image sufficient to obtain overlapping images in each spectral band. These smaller, single-band image strips, $\sim 35\times 2.6\ {\rm km}^{2}$, are mosaiced by PLC into larger images with approximately the same size as the total FOV. Different regions within the larger, mosaiced image then have different acquisition times, each corresponding to the acquisition time of the smaller, overlapping strip used in each particular region.
@@ 2.3 Time delays between positions
Figure 2: Track A in table 2.2. The circles show the locations of an aircraft identified from the individual spectral bands over a visual image (left) and over the sum of the absolute value of the seven differenced images (right). There is no circle for the NIR band because the signal-to-noise is too low to allow for precise localization. The location in the NIR band is between the two locations at the head of the track (top of the image). The spacing between the bottom two locations is longer than the others, not counting the gap due to the missing NIR location. This longer spacing is due to a shift in the mosaicing pattern that adds a time delay between exposures of the camera, $\Delta t_{camera}$, to the time delay equal to the crossing time, $\Delta t_{band}$, of one of the single-band image strips that are combined into a larger single-band image.
@@ 3 Results
Table 2: Mosaic time correction parameters $\alpha_{i}$
Track 1-5a 5-2 2-3 3-4 4-6 6-0b
A 1 0 0 0 0 0
@@
aafootnotetext: Indicates the segment measured between the positions in spectral bands 1 and 5. The segments are arranged in time order.
bbfootnotetext: Because the NIR position (band 7) is not used, $\alpha_{i}$ applies to the difference in time between the positions in bands 6 and 0.
ccfootnotetext: The ellipses indicate that a segment between the positions in two bands is not defined because one or both positions can not be determined due to low S/N. In this case the parameter $\alpha$ in the previous column applies to the segment from spectral band 3 to 6.
@@
Table 3: True velocity vs. altitude for track F
         Altitudea          true velocity
@@
aafootnotetext: Above ground level (AGL) = aircraft geo-altitude - altitude of DFW.
@@
aafootnotetext: The standard deviation is a measure of the accuracy of the velocity calculated from the PLC images, specifically the differences in the velocities of the segments of the track segments.